# On the contraction properties of Sinkhorn semigroups

### v2

@@ Appendix A The Gaussian Sinkhorn algorithm
Let us associate with some $\varpi\in{\cal S}^{+}_{d}$ the increasing map $\mbox{\rm Ricc}_{\varpi}$ from ${\cal S}^{0}_{d}$ into ${\cal S}^{+}_{d}$ defined by
$\begin{array}[]{lccl}\mbox{\rm Ricc}_{\varpi}:&{\cal S}^{0}_{d}&\mapsto&{\cal S}^{+}_{d}\\ &v&\leadsto&\mbox{\rm Ricc}_{\varpi}(v):=(I+(\varpi+v)^{-1})^{-1}\\ \end{array}.$ (101)
@@
$\varpi^{-1}:=\gamma\,\gamma^{\prime}\quad\mbox{\rm and}\quad\overline{\varpi}^{-1}:=\gamma^{\prime}\gamma\in{\cal S}^{+}_{d},\quad\mbox{\rm with}\quad\gamma:=\overline{\sigma}^{1/2}~\tau^{-1}\beta~\sigma^{1/2},$
@@
$\upsilon_{2n+1}^{-1}=I+\gamma^{\prime}~\upsilon_{2n}~\gamma\quad\mbox{and}\quad\upsilon_{2(n+1)}^{-1}=I+\gamma~\upsilon_{2n+1}~\gamma^{\prime},$ (102)
@@
$\upsilon_{2(n+1)}=\mbox{\rm Ricc}_{\varpi}\left(\upsilon_{2n}\right)\quad\mbox{and}\quad\upsilon_{2n+1}=\mbox{\rm Ricc}_{\overline{\varpi}}\left(\upsilon_{2n-1}\right),\quad\text{for $n\geq 0$}.$ (103)
@@
We also denote by $G$ a $d$-dimensional centred Gaussian random variable (r.v.) with unit covariance. The Sinkhorn transitions in (16) can be expressed as
${\cal S}_{2n}(x,dy)=\mathbb{P}\left({\cal Z}_{2n}(x)\in dy\right)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}(y,dx)=\mathbb{P}\left({\cal Z}_{2n+1}(y)\in dx\right)$
@@
${\cal Z}_{2n}(x)$ $=$ $m_{2n}+\beta_{2n}(x-m)+\tau_{2n}^{1/2}~G\quad\text{and}$
${\cal Z}_{2n+1}(y)$ $=$ $m_{2n+1}+\beta_{2n+1}(y-\overline{m})+\tau_{2n+1}^{1/2}~G.$
@@
${\cal S}^{\circ}_{2n}(y_{1},dy_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n}(y_{1})\in dy_{2}\right)\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1}(x_{1},dx_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n+1}(x_{1})\in dx_{2}\right)$
@@
${\cal Z}^{\circ}_{2n}(y)$ $:=$ $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G$
${\cal Z}^{\circ}_{2n+1}(x)$ $:=$ $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~G$
@@
From the fixed point equations (22) we have
$\overline{\sigma}=\beta^{\circ}_{2n}\overline{\sigma}\left(\beta^{\circ}_{2n}\right)^{\prime}+\tau^{\circ}_{2n}\geq\tau^{\circ}_{2n}\quad\mbox{and}\quad\sigma=\beta^{\circ}_{2n+1}\sigma\left(\beta^{\circ}_{2n+1}\right)^{\prime}+\tau^{\circ}_{2n+1}\geq\tau^{\circ}_{2n+1},$ (104)
@@
$\overline{\sigma}_{\varpi}:=\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}^{\circ}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma_{\overline{\varpi}}:=\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}^{\circ}\leq\sigma.$ (105)
@@
$\begin{array}[]{l}\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\\ \\ =(y-\overline{m})^{\prime}\left(\beta^{\circ}_{2n}\right)^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)\\ \\ =(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)^{\prime}\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)\right]\overline{\sigma}^{-1/2}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right),\end{array}$
@@
$\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]$ $=$ $(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[I-\upsilon_{2n}\right]^{2}\overline{\sigma}^{-1/2}(y-\overline{m})$ (106)
@@
The estimate in (106) ensures that
${\cal S}_{2n}^{\circ}(\psi)(y)=\mathbb{E}\left[\psi({\cal Z}^{\circ}_{2n}(y))\right]\leq(1+\lambda_{\sf min}(\varpi))^{-2}~\psi(y)+\left(\left(1-(1+\lambda_{\sf min}(\varpi))^{-2}\right)+d\right).$
@@
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})=\frac{1}{\sqrt{\mbox{\rm det}(2\pi\tau_{2n}^{\circ})}}\\ \\ \displaystyle\times~\exp{\left(-\frac{1}{2}~\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\right)}~dy_{2}\end{array}$
@@
where the second inequality follows from (105). Moreover, arguing again as above we arrive at
$\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)+2~(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\end{array}$ (107)
The inequality (107) above together with (104) readily yields
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\frac{1}{\sqrt{\mbox{\rm det}(2\pi\overline{\sigma})}}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\right)}\\ \\ \displaystyle\times~\exp{\left(-\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)\right)}~dy_{2}.\end{array}$
@@
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\varepsilon(r)~\nu_{\overline{m},\overline{\sigma}_{\varpi}/2}(dy_{2})\end{array}$
@@
This shows that the function $\psi$, and similarly the function $\varphi$ defined in (91) satisfy the drift-minorization conditions stated in (53) and (54).
Lyapunov functions in (92) and (93)
@@
$\begin{array}[]{l}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\\ \\ \leq\left[(\beta^{\circ}_{2n}(y-\overline{m}))^{\prime}\overline{\sigma}^{-1}(\beta^{\circ}_{2n}(y-\overline{m}))\right]^{1/2}+\left[(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)^{\prime}\overline{\sigma}^{-1}(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)\right]^{1/2}\\ \\ \leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2},\end{array}$
@@
$\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2}.$
This implies that the function $\psi$, and similarly the function $\varphi$ defined in (92) satisfy the drift-minorization conditions stated in (53) and (54).
The analysis of the Lyapunov functions (93) is based on the following technical lemma.
@@
$\begin{array}[]{l}\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\right)}\right]\\ \\ =\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)^{\prime}\overline{\sigma}^{-1}\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)\right]\right)}\right]\\ \\ =\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2})}}~\exp{\left(\frac{\delta}{2}~(y-\overline{m})^{\prime}~(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}~(y-\overline{m})\right)}.\end{array}$
@@
$(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}$ $=$ $(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n})^{\prime}\left(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\right)^{-1}\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}$ (108)
$\leq$ $\frac{1}{1-\delta}~(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\leq\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}~\overline{\sigma}^{-1}.$ (109)
@@
$\frac{1}{\psi(y)}~\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)$ $\leq$ $\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~I)}}$$\times$
@@
$\frac{1}{(1-\delta)(1+\lambda_{\sf min}(\varpi))^{2}}<1\quad\text{and}\quad\lim_{|y|\to\infty}\frac{1}{\psi(y)}~{\cal S}^{\circ}_{2n}(\psi)(y)=0.$
This implies that the function $\psi$, and similarly the function $\varphi$, defined in (93) satisfy the drift-minorization conditions stated in (53) and (54).

### v3

@@ Appendix A The Gaussian Sinkhorn algorithm
Let us associate with some $\varpi\in{\mathcal{S}}^{+}_{d}$ the increasing map $\mbox{\rm Ricc}_{\varpi}$ from ${\mathcal{S}}^{0}_{d}$ into ${\mathcal{S}}^{+}_{d}$ defined by
(A.1) $\begin{array}[]{lccl}\mbox{\rm Ricc}_{\varpi}:&{\mathcal{S}}^{0}_{d}&\mapsto&{\mathcal{S}}^{+}_{d}\\ &v&\leadsto&\mbox{\rm Ricc}_{\varpi}(v):=(I+(\varpi+v)^{-1})^{-1}\\ \end{array}.$
@@
$\varpi^{-1}:=\gamma\,\gamma^{\prime}\quad\mbox{\rm and}\quad\overline{\varpi}^{-1}:=\gamma^{\prime}\gamma\in{\mathcal{S}}^{+}_{d},\quad\mbox{\rm with}\quad\gamma:=\overline{\sigma}^{1/2}~\tau^{-1}\beta~\sigma^{1/2},$
@@
(A.2) $\upsilon_{2n+1}^{-1}=I+\gamma^{\prime}~\upsilon_{2n}~\gamma\quad\mbox{and}\quad\upsilon_{2(n+1)}^{-1}=I+\gamma~\upsilon_{2n+1}~\gamma^{\prime},$
@@
(A.3) $\upsilon_{2(n+1)}=\mbox{\rm Ricc}_{\varpi}\left(\upsilon_{2n}\right)\quad\mbox{and}\quad\upsilon_{2n+1}=\mbox{\rm Ricc}_{\overline{\varpi}}\left(\upsilon_{2n-1}\right),\quad\text{for $n\geq 0$}.$
@@
We also denote by $G$ a $d$-dimensional centred Gaussian random variable (r.v.) with unit covariance. The Sinkhorn transitions in (2.2) can be expressed as
${\mathcal{S}}_{2n}(x,dy)=\mathbb{P}\left({\mathcal{Z}}_{2n}(x)\in dy\right)\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}(y,dx)=\mathbb{P}\left({\mathcal{Z}}_{2n+1}(y)\in dx\right)$
@@
${\mathcal{Z}}_{2n}(x)$ $=$ $m_{2n}+\beta_{2n}(x-m)+\tau_{2n}^{1/2}~G\quad\text{and}$
${\mathcal{Z}}_{2n+1}(y)$ $=$ $m_{2n+1}+\beta_{2n+1}(y-\overline{m})+\tau_{2n+1}^{1/2}~G.$
@@
${\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2}):=\mathbb{P}\left({\mathcal{Z}}^{\circ}_{2n}(y_{1})\in dy_{2}\right)\quad\mbox{\rm and}\quad{\mathcal{S}}^{\circ}_{2n+1}(x_{1},dx_{2}):=\mathbb{P}\left({\mathcal{Z}}^{\circ}_{2n+1}(x_{1})\in dx_{2}\right)$
@@
${\mathcal{Z}}^{\circ}_{2n}(y)$ $:=$ $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G$
${\mathcal{Z}}^{\circ}_{2n+1}(x)$ $:=$ $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~G$
@@
From the fixed point equations (2.8) we have
(A.4) $\overline{\sigma}=\beta^{\circ}_{2n}\overline{\sigma}\left(\beta^{\circ}_{2n}\right)^{\prime}+\tau^{\circ}_{2n}\geq\tau^{\circ}_{2n}\quad\mbox{and}\quad\sigma=\beta^{\circ}_{2n+1}\sigma\left(\beta^{\circ}_{2n+1}\right)^{\prime}+\tau^{\circ}_{2n+1}\geq\tau^{\circ}_{2n+1},$
@@
(A.5) $\overline{\sigma}_{\varpi}:=\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}^{\circ}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma_{\overline{\varpi}}:=\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}^{\circ}\leq\sigma.$
@@
$\begin{array}[]{l}\mathbb{E}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]\\ \\ =(y-\overline{m})^{\prime}\left(\beta^{\circ}_{2n}\right)^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)\\ \\ =(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)^{\prime}\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)\right]\overline{\sigma}^{-1/2}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right),\end{array}$
@@
(A.6) $\mathbb{E}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]$ $=$ $(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[I-\upsilon_{2n}\right]^{2}\overline{\sigma}^{-1/2}(y-\overline{m})$
@@
The estimate in (A.6) ensures that
${\mathcal{S}}_{2n}^{\circ}(\psi)(y)=\mathbb{E}\left[\psi({\mathcal{Z}}^{\circ}_{2n}(y))\right]\leq(1+\lambda_{\sf min}(\varpi))^{-2}~\psi(y)+\left(\left(1-(1+\lambda_{\sf min}(\varpi))^{-2}\right)+d\right).$
@@
$\begin{array}[]{l}\displaystyle{\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2})=\frac{1}{\sqrt{\mbox{\rm det}(2\pi\tau_{2n}^{\circ})}}\\ \\ \displaystyle\times~\exp{\left(-\frac{1}{2}~\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\right)}~dy_{2}\end{array}$
@@
where the second inequality follows from (A.5). Moreover, arguing again as above we arrive at
(A.7) $\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)+2~(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right).\end{array}$
The inequality (A.7) above together with (A.4) readily yields
$\begin{array}[]{l}\displaystyle{\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2})\geq\frac{1}{\sqrt{\mbox{\rm det}(2\pi\overline{\sigma})}}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\right)}\\ \\ \displaystyle\times~\exp{\left(-\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)\right)}~dy_{2}.\end{array}$
@@
$\begin{array}[]{l}\displaystyle{\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2})\geq\varepsilon(r)~\nu_{\overline{m},\overline{\sigma}_{\varpi}/2}(dy_{2}),\end{array}$
@@
This shows that the function $\psi$, and similarly the function $\varphi$ defined in (5.10) satisfy the drift-minorization conditions stated in (4.6) and (4.7).
Lyapunov functions in (5.11) and (5.12)
@@
$\begin{array}[]{l}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\\ \\ \leq\left[(\beta^{\circ}_{2n}(y-\overline{m}))^{\prime}\overline{\sigma}^{-1}(\beta^{\circ}_{2n}(y-\overline{m}))\right]^{1/2}+\left[(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)^{\prime}\overline{\sigma}^{-1}(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)\right]^{1/2}\\ \\ \leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2},\end{array}$
@@
$\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2}.$
This implies that the function $\psi$, and similarly the function $\varphi$ defined in (5.11) satisfy the drift-minorization conditions stated in (4.6) and (4.7).
The analysis of the Lyapunov functions (5.12) is based on the following technical lemma.
@@
$\begin{array}[]{l}\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]\right)}\right]\\ \\ =\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)^{\prime}\overline{\sigma}^{-1}\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)\right]\right)}\right]\\ \\ =\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2})}}~\exp{\left(\frac{\delta}{2}~(y-\overline{m})^{\prime}~(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}~(y-\overline{m})\right)}.\end{array}$
@@
(A.8) $(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}$ $=$ $(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n})^{\prime}\left(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\right)^{-1}\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}$
(A.9) $\leq$ $\frac{1}{1-\delta}~(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\leq\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}~\overline{\sigma}^{-1}.$
@@
$\frac{1}{\psi(y)}~\mathbb{E}\left(\psi\left({\mathcal{Z}}^{\circ}_{2n}(y)\right)\right)$ $\leq$ $\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~I)}}$$\times$
@@
$\frac{1}{(1-\delta)(1+\lambda_{\sf min}(\varpi))^{2}}<1\quad\text{and}\quad\lim_{|y|\to\infty}\frac{1}{\psi(y)}~{\mathcal{S}}^{\circ}_{2n}(\psi)(y)=0.$
This implies that the function $\psi$, and similarly the function $\varphi$, defined in (5.12) satisfy the drift-minorization conditions stated in (4.6) and (4.7).